If an aqueous solution of a weak acid has a pH of 3.076 and the measured Ka for the acid is 1.4 x 10-5, then what concentration of the acid is present?
Answer #1

We have given

Ka = 1.4 \times 10-5  

pH = 3.076

We know that

pH = -log[H+]

Then

[H+] = 10-pH = 10-3.076 = 8.4  \times 10-4  

Lets see the weak acid reaction

HA(aq)  \rightleftharpoons H+(aq) + A-(aq)

The dissociation constant at equilibrium is

Ka = [H+][A-]/[HA]

At equilibrium

[H+]=[A-]

hence

Ka = [H+]2 / [HA]

[HA] = [H+]2 / Ka = ( 8.4  \times 10-4 )2 / (1.4 \times 10-5 )

= 50.4 \times 10-3 M

= 5.04 \times 10-2 M

Hence the concentration of weak acid is 5.04 \times 10-2 M